OpenStudy (anonymous):

Flying into the wind, a helicopter takes 15 minutes to travel 15 kilometers. The return flight takes 12 minutes. The wind speed remains constant during the trip. A)Find the helicopter's average speed(in kilometers per hour) for each leg of the trip.B)Write a system of linear equations that represents the situation.C) What is the helicopter's average speed in still air? What is the speed of the wind?

OpenStudy (anonymous):

average speed is distance divided by time, so going it is \(\frac{15}{15}=1\) km per minute returning it is \(\frac{15}{12}=\frac{5}{4}=1.2\) km per minute

OpenStudy (anonymous):

I need to use a system of lnear equations though.

OpenStudy (anonymous):

if you call the wind speed \(y\) and the speed in still air \(x\) then your rate going is \[x-y=1\] and your rate returning is \[x+y=1.2\]

OpenStudy (anonymous):

so the two equations are x-y=1 and x+y=1.2?

OpenStudy (anonymous):

yes, those are your two rates